# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

### v1

@@ 6 Conclusion
We introduced Ivy-Fake, the first unified, large-scale dataset for explainable AIGC detection across both images and videos, featuring over 150 000 richly annotated training samples and 18 700 evaluation examples with natural-language reasoning, and proposed Ivy Explainable Detector ( Ivy-xDetector), a vision-language architecture that jointly detects and explains synthetic content. Our model sets new state-of-the-art benchmarks in AIGC detection and explainability, and our publicly released resources provide a robust foundation for transparent, trustworthy multimodal analysis.
Limitations:Future work should optimize spatial modeling efficiency and strengthen temporal consistency. Currently, the high spatial token load (729 tokens) forces aggressive temporal downsampling, which can degrade temporal coherence and reduce accuracy in detecting subtle temporal artifacts.
@@
Appendix A Performance Evaluation of Explainability Datasets
We conduct experiments to evaluate the explanation capabilities of various models. Specifically, we assess the similarity between model-generated reasoning and reference annotations using the ROUGE-L score (Lin, 2004), which measures the longest common subsequence and reflects token-level overlap. Additionally, we adopt the LLM-as-a-judge paradigm (Zheng et al., 2023b), evaluating responses across four dimensions: Completeness, Relevance, Level of Detail, and Explanation Quality. Each response is scored using GPT-4o mini (Achiam et al., 2023) under a unified prompt that instructs the model to act as an impartial judge, assigning scores from 1 to 5. To ensure reliability, each response is rated across five independent rounds, with the final score computed as the average.
Model Image Video
Auto Metrics GPT Assisted Auto Metrics GPT Assisted
Acc F1 ROUGE-L SIM Com./Rel./Det./Exp. AVG Acc F1 ROUGE-L SIM Com./Rel./Det./Exp. AVG
Closed-source MLLMs
GPT-4o 0.766 0.759 0.155 0.683 3.69/3.79/3.67/3.80 3.74 0.828 0.813 0.150 0.682 4.02/4.13/3.99/4.12 4.07
Gemini 2.5 flash 0.668 0.656 0.223 0.709 3.29/3.33/3.27/3.33 3.30 0.787 0.784 0.188 0.678 3.85/3.93/3.82/3.92 3.88
GPT4o-mini 0.653 0.650 0.149 0.645 3.21/3.26/3.21/3.26 3.23 0.685 0.672 0.134 0.645 3.38/3.42/3.37/3.42 3.40
Open-source MLLMs
InternVL3-8B 0.614 0.605 0.159 0.680 3.04/3.07/3.04/3.07 3.05 0.632 0.616 0.165 0.629 3.13/3.16/3.12/3.16 3.14
Qwen2.5-VL-7B 0.556 0.516 0.199 0.660 2.73/2.77/2.73/2.77 2.75 0.589 0.527 0.207 0.621 2.89/2.94/2.89/2.89 2.91
Phi-3.5-Vision 0.560 0.555 0.092 0.366 2.66/2.72/2.66/2.72 2.69 0.559 0.479 0.001 0.046 2.78/2.79/2.78/2.79 2.79
Ivy-xDet 0.805 0.802 0.271 0.767 4.39/4.21/4.33/4.54 4.40 0.945 0.945 0.303 0.776 3.76/4.00/3.71/3.97 3.86
Table 5: Performance comparison of models on image and video tasks. “Auto Metrics” include Acc, F1, ROUGE-L, and SIM scores. “GPT Assisted” includes four subjective criteria: Comprehensiveness, Relevance, Detail, and Explanation, as well as their average.
As shown in Table 5, we compare our model with several leading multimodal large language models (LMMs), including Qwen2.5-7B (Bai et al., 2025), InternVL2.5-8B (Chen et al., 2024b, c), Phi-3.5-Vision, GPT-4V (Achiam et al., 2023), and Gemini 2.5 (Team et al., 2023). The results demonstrate that our approach not only achieves higher accuracy but also provides more transparent and informative explanations than all baseline models.
Appendix B Video Understanding Models and Evaluation on General Benchmarks
We further evaluate the proposed model on several video understanding benchmarks to assess its generalization capability. Specifically, we compare it against five lightweight, general-purpose video understanding models, which are VideoLLaMA3, Qwen2-VL 2B, Qwen2.5-VL-3B, InternVL2.5-2B, and InternVL3-2B, across four benchmarks, i.e., MLVU (dev), PerceptionTest, LongVideo, and VideoMME.
As shown in Table 6, our model, Ivy-Video-3B, consistently outperforms all competing methods across these benchmarks. These results highlight the strong generalization ability of our model, which, although designed for AGIC detection, achieves high accuracy across a diverse set of general-purpose video understanding tasks.
Task VideoLLaMA3 Qwen2-VL 2B Qwen2.5-VL-3B InternVL2.5-2B InternVL3-2B Ivy-Video-3B
MLVU (dev) 65.4 62.7 68.2 58.9 64.2 68.87
PerceptionTest 68 53.9 66.9 66.3 - 72.73
LongVideo 54.3 44.7 50.2 48.7 51.2 55.25
VideoMME 59.6 55.6 61.5 51.9 58.9 62.3
Table 6: Transposed performance table of various video understanding models on general video QA benchmarks.
Appendix C Effect of Incorporating Human-Annotated Labels via gemini 2.5 pro on Accuracy
@@ Appendix B Video Understanding Models and Evaluation on General Benchmarks
Appendix D Prompts
Here we provide the prompts that are mainly used in this study. As illustrated by the following figures, there are five distillation prompts distillation we used in this paper that mainly can be divided into the following three parts:
1.Prompt Template for Image Data Distillation: Since image data consists of a single frame, it can be treated as a static instance. Therefore, AIGC detection mainly focuses on identifying spatial anomalies. Detail prompt can be found in Figure 5 and  7.
2.Prompt Template for Video Data Distillation: Compared to images, video inputs provide continuous multi-frame context. This allows for detection along both spatial and temporal anomaly dimensions. etail prompt can be found in  6 and Figure 8.
3.GPT Assisted Evaluation Prompt: To assess the quality of model outputs, we design a GPT-based evaluator prompt that scores responses across four dimensions: Completeness, Relevance, Level of Detail, and Explanation. The evaluator receives a structured pair of GroundTruth and ModelOutput, each containing a <think> section (reasoning) and a <conclusion> (final judgment). The model must return a structured JSON object with integer scores (1–5) for each dimension. The prompt is provided in Figure 9.
Image Authenticity Analyst Assistant (User Prompt) Is this image real or fake? Provide the reasoning process, then give the final conclusion.
@@ Appendix D Prompts
Video Authenticity Analyst Assistant (User Prompt) Is this video real or fake? Provide the reasoning process, then give the final conclusion.
@@ 2. Spatial Analysis
- <think>: the reasoning or analytical explanation
- <conclusion>: the final judgment (e.g., real or fake)
@@ 4. Explanation
Appendix E Case Study: Qualitative Comparison of Methods
According to Figures 101112, and 13, OurMethod consistently demonstrates superior performance in detecting both spatial and temporal anomalies, providing stronger generalization and robustness compared to existing baselines.
@@ Appendix E Case Study: Qualitative Comparison of Methods
Figure 12: Video example 1 demonstrating that Ivy-xDetector effectively captures temporal inconsistencies overlooked by baseline models.

### v2

@@ 5 Conclusion
We introduced Ivy-Fake, the first unified, large-scale dataset for explainable AIGC detection across both images and videos, featuring over 131K richly annotated training samples and 5,000 evaluation examples with natural-language reasoning, and proposed Ivy Explainable Detector ( Ivy-xDetector), a MLLM that jointly detects and explains synthetic content. Our model sets superior benchmarks in AIGC detection and explainability, and our publicly released resources provide a robust foundation for transparent, trustworthy multimodal analysis. We believe that such fine-grained explainability data Ivy-Fake can catalyze new research directions for the AIGC detection community, enabling deeper understanding of synthetic content and more principled model development.
Limitations: Future work will concentrate on tailoring the benchmark toward domain-specific scenarios, including explainability-focused datasets for DeepFake and document manipulation analysis. Nevertheless, our experiments suggest that the proposed approach exhibits strong scalability and generalization capabilities, indicating the potential for even better performance when applied to larger models such as 32B or 72B.
@@
[…]
Supplementary Material
6 License
For all datasets in our experiments and datasets, FLUX.1 [labs2025flux1kontextflowmatching] and Lora Flux [labs2025flux1kontextflowmatching] falls under the FLUX.1 [dev] Non-Commercial License.WildFake [wildfake_2025] is released under an open-source license, making it freely available for research and non-commercial use; Qwen-Image [wu2025qwenimagetechnicalreport] is licensed under Apache 2.0; Sora [sora_videoworldsimulators2024] is a commercial, non-open-source text-to-video model provided via ChatGPT Plus/Pro; user-generated content is owned by the user and allows (subject to usage policies) non-commercial use, but the model itself is closed-source.
Stable Diffusion [svd_blattmann2023stable], while originally open-source under Creative ML OpenRAIL-M (e.g., versions 1.X, 2.1, SDXL), has transitioned in newer versions (SD3.x, SD3.5) to the Stability AI Community License, which permits free use for individuals or entities with annual revenue under USD 1 million, but requires a paid Enterprise License for larger-scale or commercial usage.
During Ivy-Fake construction and experimentation, we accessed Gemini API and ChatGPT API to ensure annotation and evaluation’s quality and reproducibility. We explicitly state that all data generated using these commercial APIs are used solely for academic research and non-commercial purposes, fully complying with the respective API usage agreements and ethical guidelines. No Gemini- or ChatGPT-derived content is redistributed for any commercial training, deployment, or monetization purposes.
7 Related Work
7.1 Methods for Synthetic Content Detection
Due to growing concerns about the misuse of synthetic data [deng2024survey], research on AI-generated content (AIGC) detection has expanded rapidly in recent years. Most existing models for AI-generated images and videos formulate the task as binary classification, simply predicting whether the content is ”real” or ”fake.” Representative examples include CNN-based AIGVDet [aigvdet_bai2024ai], CNNSpot [cnnspot_wang2020cnn] and Transformer-based models such as DIRE [dire_wang2023dire] and AIDE [aide_yan2025a]. Meanwhile, several works have explored the application of multimodal large language models (MLLMs) to AIGC detection, including Synartifact [cao2024synartifact] and Bi-LORA [keita2025bi]. However, these approaches largely overlook the importance of interpretability in AIGC detection.
Some efforts attempt to introduce interpretability by leveraging spatial annotations [dong2022explaining] or frequency-domain artifact analysis [zhang2023perceptual]. Nevertheless, the resulting explanations are often difficult for humans to comprehend, as they lack clarity in natural language. This limitation is particularly evident in the video domain, where AI-generated content frequently exhibits obvious flaws, e.g., incoherent frame transitions and object inconsistency, that are easily noticed and reasoned about by humans [deng2024survey]. FakeClue [fakeclue_wen2025spot] introduces the use of vision-language models (VLMs) to provide interpretability for image-level detection, but it does not offer a unified framework that integrates both images and videos.
7.2 Datasets for Synthetic Content Detection.
Early datasets for synthetic content detection, such as CNNSpot [cnnspot_wang2020cnn], primarily collected fake images generated by GAN-based models [gan1_goodfellow2014generative, gan2_zhu2017unpaired, gan3_brock2018large]. However, with the advent of more advanced generative architectures like diffusion models [diff1_ho2020denoising, diff2_dhariwal2021diffusion, diff3_rombach2022high, diff4_hertz2022prompt, diff5_nichol2021glide] and their variants, the authenticity of generated content has significantly increased, making it more challenging for detection models to discern. This has spurred the development of newer datasets, including ArtiFact [cao2024synartifact], GenImage [genimage_2023], and WildFake [wildfake_2025]. GenImage [genimage_2023] comprises images from the 1000 ImageNet [imagenet_ILSVRC15] categories, generated by eight state-of-the-art generators such as Stable Diffusion [diff3_rombach2022high] and Midjourney. Nevertheless, these datasets predominantly focus on image-based content. More recently, datasets emphasizing interpretability have also been introduced. FakeClue [fakeclue_wen2025spot] contains a large amount of image data with explainability annotations but lacks video data. LOKI [loki_ye2024loki] offers data across 26 different categories and includes 18,000 distinct questions; however, its volume of multimodal data is relatively small and primarily suited for evaluation rather than comprehensive model training. Therefore, a critical gap exists for a unified benchmark encompassing both image and video modalities to rigorously evaluate the performance of contemporary AIGC detectors.
Figure 6: Final Reinforcement Learning Training Curves.
8 GRPO
Following DeepSeek-R1 [deepseekr1_guo2025deepseek], we adapt the Group Relative Policy Optimization (GRPO) [shao2024deepseekmathpushinglimitsmathematical], an online RL algorithm designed to maximize the advantage of generated completions while constraining policy divergence from a reference model. We formalize our training process of Ivy-xDetectorusing GRPO below. Let $p$ denote a sampled prompt, and let $o_{1},o_{2},...,o_{n}$} be a group of completions generated by the current policy $\pi_{\theta}$. For each completion $G_{i}$, a reward $r_{i}$ is computed using a rule-based reward function. The group advantage for each completion is then calculated as:
$\hat{A}_{i,t}=\frac{r_{i}-mean(r)}{std(r)}$ (1)
$\mathcal{L}_{\mathrm{GRPO}}(\theta)=-\frac{1}{K}\sum_{k=1}^{K}A^{(k)}\,\ell^{(k)}(\theta)\;+\;\beta\,\mathrm{KL}\!\left(\pi_{\theta}(\cdot|x)\,\|\,\pi_{\mathrm{ref}}(\cdot|x)\right).$ (2)
where $\beta$ is a coefficient that balances advantage maximization and KL regularization, and the clipping operator $clip(\dots,1-\epsilon,1+\epsilon)$ constrains the update magnitude. To regularize policy updates, we estimate the token-level Kullback-Leibler (KL) divergence between the current policy $\pi_{\theta}$ and a reference policy $\pi_{ref}$.
$\mathbb{D}_{\text{KL}}\big[\pi_{\theta}\big\|\pi_{\text{ref}}\big]=\frac{\pi_{\text{ref}}(o_{i,t}\mid p,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid p,o_{i,<t})}-\log\frac{\pi_{\text{ref}}(o_{i,t}\mid p,o_{i,<t})}{\pi_{\theta}(o_{i,t}\mid p,o_{i,<t})}-1.$ (3)
Reward Model. For effective RL, we employ a rule-based reward that consists of accuracy and format rewards. We introduce a solid accuracy reward for AIGC Detection, which utilizes distinct functions to evaluate binary classification task. This allows for a more appropriate assessment based on the expected answer type.
  • Accuracy Reward: The accuracy reward assigns a score of 1 if the label in <conclusion> exactly matches the ground-truth classification real and fake and 0 otherwise.
  • Format Reward: The format reward assigns a score of 1 if the output strictly follows the structural requirements by enclosing the reasoning within <think></think> tags and the final decision within <conclusion></conclusion> tags, and 0 otherwise.
9 Effect of Incorporating Human-Annotated Labels via gemini 2.5 pro on Accuracy
@@ 8 GRPO
10 Data Distribution
11 Additional Experiment
11.1 Synthetic Image Detection
We evaluated our AIGC detector on the GenImage [genimage_2023] and Chameleon [aide_yan2025a] benchmarks.
As shown in Table 5, GenImage comprises seven subsets generated by leading models, i.e., Midjourney, Stable Diffusion v1.4 & v1.5, ADM, GLIDE, Wukong, VQDM, and BigGAN. We compared against five state-of-the-art detectors, which are CNNSpot [cnnspot_wang2020cnn], F3Net [qian2020thinking], DIRE [dire_wang2023dire], GenDet [zhu2023gendet], PatchCraft [aigcdetectionbenchmark_zhong2023patchcraft], and AIDE [aide_yan2025a].
Notably, our detector surpasses previous state-of-the-art methods such as AIDE (86.88%) and PatchCraft (82.30%) by a large margin, even though it is trained under a unified multimodal paradigm rather than a generator-specific setting. The improvement is especially pronounced on challenging generators like ADM, GLIDE, and BigGAN, where conventional CNN- or patch-based detectors often fail to capture high-frequency inconsistencies or generalized texture patterns.
Notably, our model is trained with only 130K image samples—fewer than GenImage’s 2M+ training images—yet still achieves superior generalization performance.
Image Authenticity Analyst Assistant (User Prompt) Is this image real or fake?
@@ 11.1 Synthetic Image Detection
Video Authenticity Analyst Assistant (User Prompt) Is this video real or fake?
@@ 2. Spatial Analysis
11.2 Synthetic Video Detection
As shown in Table 6, we evaluate the performance of our detector on the GenVideo benchmark [demamba].
Specifically, Ivy-xDetector achieves an average F1 score of 95.26% and recall of 95.28%, substantially surpassing the prior state-of-the-art DeMamba (F1 = 90.20%). The improvement is particularly evident on challenging categories such as HotShot, Lavie, and Show-1, where previous methods tend to overfit to specific generative distributions or temporal artifacts. Our results indicate that the unified MLLM-based detection paradigm not only captures spatial inconsistencies but also learns transferable temporal-spatial representations, leading to more robust generalization across unseen generative models.
[…]
12 Prompts
Default System Prompt You are an AI-generated content detector. Given a single media (image or video), classify it as real or fake. Provide detailed reasoning inside the <think></think> tags, including your step-by-step thought process.
Your output must begin with <think>
n and end with </conclusion>.
Then output exactly one word in lowercase—either real or fake—wrapped in <conclusion></conclusion>.
Do not include any other words. If uncertain, choose the most likely class.
Figure 11: Default System Prompt for Training and Evaluation
Here we provide the prompts that are mainly used in this study. The default system prompt can be seen in Figure11. As illustrated by the following figures, there are five distillation prompts distillation we used in this paper that mainly can be divided into the following three parts:
Prompt Template for Image Data Distillation: Since image data consists of a single frame, it can be treated as a static instance. Therefore, AIGC detection mainly focuses on identifying spatial anomalies. Detail prompt can be found in Figure 7 and  9.
Prompt Template for Video Data Distillation: Compared to images, video inputs provide continuous multi-frame context. This allows for detection along both spatial and temporal anomaly dimensions. Dtail prompt can be found in Figure 8 and Figure 10.
GPT Assisted Evaluation Prompt: To assess the quality of model outputs, we design a GPT-based evaluator prompt that scores responses across four dimensions: Completeness, Relevance, Level of Detail, and Explanation. The evaluator receives a structured pair of GroundTruth and ModelOutput, each containing a <think> section (reasoning) and a <conclusion> (final judgment). The model must return a structured JSON object with integer scores (1–5) for each dimension. The prompt is provided in Figure 12.
@@ 12 Prompts
- ¡think¿: the reasoning or analytical explanation
- ¡conclusion¿: the final judgment (e.g., real or fake)
@@ 4. Explanation
13 Training Detail
The RL training process of Ivy-xDetector is illustrated in Fig.6. Our GRPO training exhibits a stable increase in reward, eventually reaching 0.95. However, we also observe that both the clip ratio and the maximum response length surge sharply around 200–300 steps, suggesting that a portion of the training samples may be overly challenging for the model.
14 Case Study: Qualitative Comparison of Methods
According to Figures 131415, and 16, Ivy-xDetector consistently demonstrates superior performance in detecting both spatial and temporal anomalies, providing stronger generalization and robustness compared to existing baselines.